# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ Results.
Figure 1: The zeros of several instances of polynomial tau functions (all for ${K}=0$).
In the figure an almond-shaped region, which we call “Eye of the Tiger” ($EoT$) is visible. We will show that for fixed values of $K,\rho$ and as $n$ tends to infinity, the zeros of $\mathcal{T}_{n}(s)$ will eventually vacate any closed set $\mathcal{K}$ not intersecting the (closure of) $EoT$ or, which is the same, the zeros will eventually lie in an $\epsilon$–fattening of $EoT$.
This phenomenon is well illustrated by Fig. 3, top right pane; indeed we can see that certain zeros lie outside the region (for the specific value of $n=17$). If we were to plot the zeros for the same values of the parameters ${K},{\rho}$, but for larger values of $n$, the zeros would gradually drift towards the region $EoT$. The method of proof does not allow us to conclusively show that all the zeros will eventually fall in the interior of $EoT$. See Theorem 5.2.
The boundary of the region $EoT$ (marked in black) is determined by the implicit equation in the $s$–plane
$\ln\left|\frac{2}{s}+\sqrt{\frac{4}{s^{2}}+1}\right|-\mathrm{Re}\,\left(\sqrt{1+\frac{s^{2}}{4}}\right)=0,$ (1.18)
which follows from (4.20), (4.26) below. Within the set $EoT$ the zeros of $\mathcal{T}_{n}(s)$ (1.17) arrange themselves in a semi-regular lattice. We give a description of this lattice in terms of a “quantization condition” described in Section 5.2.4, in particular (5.2.1) (for $\mathrm{Re}\,(s)>0$) or (5.2.2) (for $\mathrm{Re}\,(s)<0$). The two quantization conditions (5.2.1), (5.2.2) determine an asymptotic grid (shown in Fig. 3 in green and blue thin lines) at whose vertices the zeros are approximately located. We do not estimate rigorously the rate of convergence, but the numerical evidence is quite striking even for relatively small values of $n$.
Remark 1.2
Our analysis assumes that $\rho$ is not integer. This is an essential assumption, not of technical nature. It is sufficient to show a few plots of the zeros of $\tau_{n}(t;K,\rho)$ to see that the pattern of zeros inside $EoT$ has completely collapsed. For example, it is easy to see that if $\rho=0$ the tau function reduces to a monomial. On the technical level, for the asymptotic analysis in the interior of $EoT$ we need the density of our measure to have a branch-cut discontinuity from $z=0$ to $z=1$. A few plots with integer $\rho$ are indicated in Fig. 2 to illustrate the point.
@@
Figure 2: Some plot of the zeros of the tau function for $(n,\rho)=(30,1),(30,2),(30,3)$ ($K=0$ in all cases). Observe that for $\rho$ integer our description of the distribution of zeros does not apply and in particular the region $EoT$ remains largely empty, with an “eyelash” effect. There is a high multiplicity zero at the origin. As $n$ increases but $\rho$ remains fixed, the “eyelashes” become thinner around the edge of $EoT$. A similar phenomenon was observed, for rational solutions of PIII, in [13]. We are not going to discuss the case $\rho\in{\mathbb{Z}}$ in this paper.
2 Semiclassical orthogonal polynomials and tau functions of Painlevé type
@@
Definition 2.1
Given a pair of polynomials $(A,B)$, with $B$ monic, a semiclassical moment functional of type $(A,B)$ is a linear map ${\mathcal{M}}:{\mathbb{C}}[z]\to{\mathbb{C}}$ on the space of polynomials in the indeterminate variable $z$ such that for all $p\in{\mathbb{C}}[z]$
@@
${\mathcal{M}}[p_{n}(z)z^{j}]=0,\ \ \ j=0,1,\dots,n-1.$ (2.2)
Given a moment functional ${\mathcal{M}}$ its moments are the (in general complex) numbers
$\mu_{j}:=\mathcal{M}[z^{j}].$ (2.3)
It is possible to show that these moment functionals admit an integral representation as follows: define the symbol [^3] We adopt the terminology common in the literature on Toeplitz operators. as the function $\theta$ such that
$\theta^{\prime}(z):=-\frac{B^{\prime}(z)+A(z)}{B(z)}.$ (2.4)
@@
${\mathcal{M}}\left[p(z)\right]=\int_{\gamma}p(z){\rm e}^{\theta(z)}{\mathrm{d}}z.$ (2.5)
@@
Indeed it was shown in [7] that any deformation of the coefficients of $A,B$ that preserves all the residues of $\theta^{\prime}$ is an isomonodromic deformation for a suitably defined differential equation in the complex plane of rank $2$. It was determined by the Japanese school several decades ago [32, 33] that to any such isomonodromic deformation we can associate a tau function, namely a function of the isomonodromic deformation parameters (i.e. of the coefficients of $A,B$ in this case). While this is not an appropriate venue to review all applications of tau functions, suffice to say here (see Rem. 5.3 of [7]) that by appropriate choices of the symbol and corresponding contours of integration $\gamma$ we can construct tau function for all the Painlevé equations II–VI. We will make below one such choice which has the additional property of producing polynomial moments, and hence polynomial tau functions.
@@ 2.1 Special case of semiclassical functional of PV type
We specialize the previous description to the case of the following symbol depending on the parameters $t,{{K}},{{\rho}}$
${\theta(z)}:=\frac{t}{z}+{{\rho}}\ln\left(1-\frac{1}{z}\right)+{{K}}\ln z,\ \ \ {\rm e}^{\theta(z)}=\left(1-\frac{1}{z}\right)^{{\rho}}z^{{{K}}}{\rm e}^{\frac{t}{z}}.$ (2.6)
In (2.6) all the logarithms are principal and $\theta$ is analytic in ${\mathbb{C}}\setminus(-\infty,1]$. Since ${K},\ {\rho}$ are the residues of $\theta^{\prime}{\mathrm{d}}z$, the only isomonodromic parameter is $t$ in the above expression. Note that if ${{K}}\in{\mathbb{Z}}$ (which is going to be our main focus shortly) the weight function ${\rm e}^{\theta(z)}$ is analytic in ${\mathbb{C}}\setminus[0,1]$. Furthermore we will avail ourselves of the freedom of deforming the segment $[0,1]$ to a smooth arc with the same endpoints when needed. The corresponding moment functionals, in the terminology of Def. 2.1 above, are of type
$(A,B)=\bigg{(}{{K}}z^{2}+({{\rho}}+{{K}}+t)z+t,\ z^{2}(z-1)\bigg{)}.$ (2.7)
@@
There are some special instances, depending on the values of the parameters ${{K}},{{\rho}}$; if ${{\rho}}=0,1,2,\dots,$ then we can replace $\widetilde{\gamma}$ with a contour terminating at $z=1$.
Most important for our consideration is the case ${K}\in{\mathbb{Z}},{{\rho}}\not\in{\mathbb{Z}}$; in these cases, we can replace $\gamma$ with the homotopy class of a circle of radius $R>1$ in ${\mathbb{C}}\setminus[0,1]$.
The latter case is relevant for the construction of rational solutions of the fifth Painlevé equation: indeed, if ${K}\in{\mathbb{Z}}$ the moments are polynomials in the variable $t$
$\mu_{j}(t)=\oint_{|z|=R}z^{j}{\rm e}^{\theta(z)}{\mathrm{d}}z,\ \ \ \deg\mu_{j}(t)=j+{K}+1.$ (2.8)
In our paper we consider the Hankel determinants:
$\tau_{n}(t;{K},{\rho}):=\det\bigg{[}\mu_{a+b-2}(t)\bigg{]}_{a,b=1}^{n}.$ (2.9)
They are also polynomials of degree
$\deg\tau_{n}(t)=n(n+{K}).$ (2.10)
In fact the determinant of leading terms in $t$ of the matrix of moments gives
$\tau_{n}(t;{K},{\rho})=t^{n(n+{K})}\det\left[\frac{1}{({K}+a+b-1)!}\right]_{a,b=1}^{n}\left(1+\mathcal{O}(t^{-1})\right).$ (2.11)
It is simple to establish that the leading coefficient does not vanish, so that the degree of $\tau_{n}$ is exactly $n(n+{K})$. Indeed, one may notice that the determinant is the Wronskian, evaluated at $x=1$ and up to reversal of order of rows, of the monomials $f_{j}(x)=\frac{x^{{K}+2n-j}}{({K}+2n-j)!}$, $j=1,\dots,n$. This Wronskian is itself a monomial, and hence it does not vanish at $x=1$.
Remark 2.1 (Comparison with [23])
To connect the moments with known functions, we perform the change of variable $\frac{1}{z}=\frac{s}{s-1}$ in the defining integral of the moments $\mu_{j}(t)$ (2.8) ; after a simple computation the expression for the moment becomes
$\mu_{j}(t)=(-1)^{j+{K}}\oint_{|s|=r}\frac{{\rm e}^{t\frac{s}{s-1}}}{(1-s)^{\nu+1}s^{n+1}}\frac{{\mathrm{d}}s}{2i\pi},\qquad\nu={\rho}-j-{K}-1,\ \ n=j+{K}+1.$ (2.12)
@@
[…]
from which we conclude
$\mu_{j}(t)=(-1)^{j+{K}}L^{{\rho}-j-{K}-1}_{j+{K}+1}(t).$ (2.14)
Using finally the simple relation $\partial_{t}^{k}\mu_{j}(t)=\mu_{j-k}(t)$, we can rewrite the Hankel determinant (2.9) as
$\tau_{n}(t;{K},{\rho})=(-1)^{{K}n}\det\left[\left(\frac{{\mathrm{d}}}{{\mathrm{d}}t}\right)^{a+b}\mu_{2n-2}(t)\right]_{a,b=0}^{n-1}=(-1)^{{K}n}\det\left[\left(\frac{{\mathrm{d}}}{{\mathrm{d}}t}\right)^{a+b}L^{({\rho}-{K}-2n+1)}_{2n-1+{K}}(t)\right]_{a,b=0}^{n-1}.$ (2.15)
With reference to the notation of [23], Definition 3.1 ibidem, the relationship is then
$\tau_{n}(t;{K},{\rho})=(-1)^{{K}n}T^{({\rho}-{K}-2n)}_{n+{K}-1,n}(t)=(-1)^{{K}n}t^{\frac{n(1-n)}{2}}a_{n+{K}-1,n}\,\tau^{({\rho}-{K}-2n)}_{n+{K}-1,n}(t)$ (2.16)
where, in the last identity we have used Lemma 4.5 of loc. cit. to connect with the definition (1.2). The proportionality constant $a_{m,n}$ is defined in the same Lemma. $\triangle$
@@
Remark 2.2
We are considering the moment functional of type (2.7) that corresponds to the choice of contour $\gamma$ surrounding the segment $[0,1]$ in the complex plane. The general theory of moment functionals discussed earlier allows for other choices. While these do not lead to polynomial moments (in the variable $t$), their Hankel determinants would, by the same reasoning in [7], provide isomonodromic tau functions of Painlevé V. Up to the map $z\mapsto 1/z$ and setting ${\rho}={K}=0$ and choosing the contour of integration as the segment $[0,1]$ (the so–called “hard-edge case”), one would recover the moment functional appearing in [5]. In that case the orthogonal polynomials are the so-called “kissing polynomials”. It would be possible to perform a similar asymptotic analysis of the corresponding tau functions for large $n$ and determine the zero distribution, despite the tau functions not being polynomials. $\triangle$
@@ 2.2 From orthogonal polynomials to Lax Pairs
Problem 2.1 (RHPY)
Given $n\in\mathbb{N}$, find a $2\times 2$ matrix–valued function $Y(z)=Y_{n}(z)$ such that
  1. 1.
@@
    $Y(z_{+})=Y(z_{-})\begin{bmatrix}1&{\rm e}^{\theta(z)}\\ 0&1\end{bmatrix},\qquad\forall z\in\gamma$ (2.17)
@@
  3. 3.
    As $z\to\infty$ the matrix $Y_{n}(z)$ admits an asymptotic expansion of the form
    $Y_{n}(z)=\left(\mathbf{1}+\sum_{\ell=1}^{\infty}\frac{Y_{n}^{(\ell)}}{z^{\ell}}\right)z^{n\sigma_{3}},\ \ \ \ \sigma_{3}:=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},$ (2.18)
    where the coefficient matrices $Y_{n}^{(\ell)}$ are independent of $z$ and $\mathbf{1}$ denotes the identity matrix of the appropriate size (here, $2\times 2$).
In the following theorem we condense the essential results that are at the core of much of the theory of orthogonal polynomials. The theorem is mostly the reformulation of results of [28] together with notational adaptations to the current context.
Theorem 2.1 ([28])
Consider the RHP 2.1;
  1. 1.
    If the solution exists it is unique and $\det Y_{n}(z)\equiv 1$.
  2. 2.

### v3

@@ 4. $(IV)$
Figure 1: The zeros of several instances of polynomial tau functions (all for K=0{K}=0).
In the figure, an almond-shaped region, which we call “Eye of the Tiger” ($\rm EoT$}) is visible. We will show that for fixed values of $K$, $\rho$ and as $n$ tends to infinity, the zeros of $\mathcal{T}_{n}(s)$ will eventually vacate any closed set $\mathcal{K}$ not intersecting the (closure of) $\rm EoT$} or, which is the same, the zeros will eventually lie within $\epsilon$-distance of $\rm EoT$}.
Namely, for every $K\in{\mathbb{N}}$, $\rho\in{\mathbb{C}}\setminus{\mathbb{Z}}$, $\epsilon>0$, there is an $N_{0}=N_{0}(K,\rho,\epsilon)$ such that all the zeros of $\mathcal{T}_{n}(s)$ are inside the $\epsilon$-fattening ${\rm EoT}^{\epsilon}$. We recall that an $\epsilon$-fattening of a bounded set $X\subset{\mathbb{C}}$} is the (open) set $X^{\epsilon}:=\bigcup_{z\in X}\mathbb{D}_{z}(\epsilon)$, $\mathbb{D}_{z}(\epsilon):=\{w\mid|w-z|<\epsilon\}$. We do not provide an estimate of the threshold value $N_{0}$.
This phenomenon is well illustrated by Figure 3, top right pane. Indeed, we can see that certain zeros lie outside the region (for the specific value of $n=17$). If we were to plot the zeros for the same values of the parameters K{K}, $\rho$}, but for larger values of $n$, the zeros would gradually drift towards the region $\rm EoT$}. The method of proof does not allow us to conclusively show that all the zeros will eventually fall in the interior of $\rm EoT$}. See Theorem 5.8.
The boundary of the region $\rm EoT$} (marked in black) is determined by the implicit equation in the $s$-plane
$\ln\left|\frac{2}{s}+\sqrt{\frac{4}{s^{2}}+1}\right|-\operatorname{Re}\left(\frac{s}{2}\sqrt{\frac{4}{s^{2}}+1}\right)=0,$ (1.6)
which follows from (4.5), (4.7) below. Note that in the above expression the square roots need to be understood with branch-cuts extending from $2{\rm i}$ vertically upwards, and from $2{\rm i}$ vertically downwards. With this choice the function on the left side of (1.6) is continuous in ${\mathbb{C}}\setminus\{0\}$} and even.
Within the set $\rm EoT$}, the zeros of $\mathcal{T}_{n}(s)$ (1.5) arrange themselves in a semi-regular lattice. We give a description of this lattice in terms of a “quantization condition” described in Section 5.2.4, in particular (5.41) (for $\operatorname{Re}(s)>0$) or (5.47) (for $\operatorname{Re}(s)<0$). The two quantization conditions (5.41), (5.47) determine an asymptotic grid (shown in Figure 3 in green and blue thin lines) at whose vertices the zeros are approximately located. We do not estimate rigorously the rate of convergence, but the numerical evidence is quite striking even for relatively small values of $n$.
Remark 1.3.
Our analysis assumes that $\rho$ is not integer. This is an essential assumption, not of technical nature. It is sufficient to show a few plots of the zeros of $\mathcal{T}_{n}(s)$ (1.5) to see that the pattern of zeros inside $\rm EoT$} has completely collapsed. For example, it is easy to see that if $\rho=0$ the tau function reduces to a monomial. On the technical level, for the asymptotic analysis in the interior of $\rm EoT$} we need the density of our measure to have a branch-cut discontinuity from z=0{z=0} to $z=1$. A few plots with integer $\rho$ are indicated in Figure 2 to illustrate the change of behaviour.
@@
Figure 2: Some plot of the zeros of the tau function $\mathcal{T}_{n}(s)$ (1.5) for $(n,\rho)=(30,1),(30,2),(30,3)$ ($K=0$ in all cases). Observe that for $\rho$ integer our description of the distribution of zeros does not apply and in particular the region $\rm EoT$} remains largely empty, with an “eyelash” effect. There is a high multiplicity zero at the origin. As $n$ increases but $\rho$ remains fixed, the “eyelashes” become thinner around the edge of $\rm EoT$}. A similar phenomenon was observed, for rational solutions of PIII, in [14]. We are not going to discuss the case $\rho\in{\mathbb{Z}}$ in this paper. Clearly the behaviour of the zeros undergoes a substantial change, as they appear to be accumulating along part of the boundary of $\rm EoT$}.
2 Semiclassical orthogonal polynomials and tau functions
of Painlevé type
@@ 2 Semiclassical orthogonal polynomials and tau functions
Definition 2.1.
Given a pair of polynomials $(A,B)$, with $B$ monic, a semiclassical moment functional of type $(A,B)$ is a linear map ${\mathcal{M}}\colon{\mathbb{C}}[z]\to{\mathbb{C}}$ on the space of polynomials in the indeterminate variable $z$ such that for all $p\in{\mathbb{C}}[z]$
@@
${\mathcal{M}}\big[p_{n}(z)z^{j}\big]=0,\qquad j=0,1,\dots,n-1.$
Given a moment functional $\mathcal{M}$} its moments are the (in general complex) numbers $\mu_{j}:=\mathcal{M}\big[z^{j}\big]$. It is possible to show that these moment functionals admit an integral representation as follows: define the symbol [^1] We adopt the terminology common in the literature on Toeplitz operators. as the function $\theta$ such that
$\theta^{\prime}(z):=-\frac{B^{\prime}(z)+A(z)}{B(z)}.$
@@
${\mathcal{M}}\left[p(z)\right]=\int_{\gamma}p(z){\rm e}^{\theta(z)}{\mathrm{d}}z.$
@@
Indeed, it was shown in [7] that any deformation of the coefficients of $A$, $B$ that preserves all the residues of $\theta^{\prime}$ is an isomonodromic deformation for a suitably defined differential equation in the complex plane of rank $2$. It was determined by the Japanese school several decades ago [32, 33] that to any such isomonodromic deformation we can associate a tau function, namely a function of the isomonodromic deformation parameters (i.e., of the coefficients of $A$, $B$ in this case). While this is not an appropriate venue to review all applications of tau functions, suffice to say here (see [7, Remark 5.3]) that by appropriate choices of the symbol and corresponding contours of integration $\gamma$ we can construct tau function for all the Painlevé equations II–VI. We will make below one such choice which has the additional property of producing polynomial moments, and hence polynomial tau functions.
@@ 2.1 Special case of semiclassical functional of PV type
We specialize the previous description to the case of the following symbol depending on the parameters $t$, K{{K}}, $\rho$}}:
${\theta(z)}:=\frac{t}{z}+{{\rho}}\ln\left(1-\frac{1}{z}\right)+{{K}}\ln z,\qquad{\rm e}^{\theta(z)}=\left(1-\frac{1}{z}\right)^{{\rho}}z^{{{K}}}{\rm e}^{\frac{t}{z}}.$ (2.2)
In (2.2), all the logarithms are principal and $\theta$ is analytic in ${\mathbb{C}}\setminus(-\infty,1]$. Since K{K}, $\rho$} are the residues of $\theta^{\prime}{\mathrm{d}}z$, the only isomonodromic parameter is $t$ in the above expression. Note that if ${{K}}\in{\mathbb{Z}}$ (which is going to be our main focus shortly) the weight function ${\rm e}^{\theta(z)}$ is analytic in ${\mathbb{C}}\setminus[0,1]$}. Furthermore, we will avail ourselves of the freedom of deforming the segment $[0,1]$ to a smooth arc with the same endpoints when needed. The corresponding moment functionals, in the terminology of Definition 2.1 above, are of type
$(A,B)=\bigl({{K}}z^{2}+({{\rho}}+{{K}}+t)z+t,z^{2}(z-1)\bigr).$ (2.3)
@@
There are some special instances, depending on the values of the parameters ${{K}},{{\rho}}$. If ${{\rho}}=0,1,2,\dots$, then we can replace $\widetilde{\gamma}$ with a contour terminating at $z=1$.
Most important for our consideration is the case ${K}\in{\mathbb{Z}}$, $\rho$$\not\in{\mathbb{Z}}$. In these cases, we can replace $\gamma$ with the homotopy class of a circle of radius $R>1$ in ${\mathbb{C}}\setminus[0,1]$.
The latter case is relevant for the construction of rational solutions of the fifth Painlevé equation. Indeed, if ${K}\in{\mathbb{Z}}$ the moments are polynomials in the variable $t$,
$\mu_{j}(t)=\oint_{|z|=R}z^{j}{\rm e}^{\theta(z)}{\mathrm{d}}z,\qquad\deg\mu_{j}(t)=j+{K}+1.$ (2.4)
In our paper, we consider the Hankel determinants
$\tau_{n}(t;{K},{\rho}):=\det\bigl[\mu_{a+b-2}(t)\bigr]_{a,b=1}^{n}.$ (2.5)
They are also polynomials of degree $\deg\tau_{n}(t)=n(n+{K})$. In fact, the determinant of leading terms in $t$ of the matrix of moments gives
$\tau_{n}(t;{K},{\rho})=t^{n(n+{K})}\det\left[\frac{1}{({K}+a+b-1)!}\right]_{a,b=1}^{n}\left(1+\mathcal{O}\bigl(t^{-1}\bigr)\right).$
It is simple to establish that the leading coefficient does not vanish, so that the degree of $\tau_{n}$ is exactly n(n+K)n(n+{K}). Indeed, one may notice that the determinant is the Wronskian, evaluated at x=1{x=1} and up to reversal of order of rows, of the monomials $f_{j}(x)=\frac{x^{{K}+2n-j}}{({K}+2n-j)!}$, $j=1,\dots,n$. This Wronskian is itself a monomial, and hence it does not vanish at $x=1$.
Remark 2.2 (comparison with [23]).
To connect the moments with known functions, we perform the change of variable $\frac{1}{z}=\frac{s}{s-1}$ in the defining integral of the moments $\mu_{j}(t)$ (2.4). After a simple computation, the expression for the moments becomes
$\mu_{j}(t)=(-1)^{j+{K}}\oint_{|s|=r}\frac{{\rm e}^{t\frac{s}{s-1}}}{(1-s)^{\nu+1}s^{n+1}}\frac{{\mathrm{d}}s}{2{\rm i}\pi},\qquad\nu={\rho}-j-{K}-1,\qquad n=j+{K}+1.$
@@
$L^{(\nu)}_{n}(t)=\oint_{|s|=r}\frac{{\rm e}^{t\frac{s}{s-1}}}{(1-s)^{\nu+1}s^{n+1}}\frac{{\mathrm{d}}s}{2{\rm i}\pi},\qquad 0<r<1,$
from which we conclude $\mu_{j}(t)=(-1)^{j+{K}}L^{{\rho}-j-{K}-1}_{j+{K}+1}(t)$. Using finally the simple relation $\partial_{t}^{k}\mu_{j}(t)\allowbreak=\mu_{j-k}(t)$, we can rewrite the Hankel determinant (2.5) as
$\tau_{n}(t;{K},{\rho})={}$ $(-1)^{{K}n}\det\left[\left(\frac{{\mathrm{d}}}{{\mathrm{d}}t}\right)^{a+b}\mu_{2n-2}(t)\right]_{a,b=0}^{n-1}$
$={}$ $(-1)^{{K}n}\det\left[\left(\frac{{\mathrm{d}}}{{\mathrm{d}}t}\right)^{a+b}L^{({\rho}-{K}-2n+1)}_{2n-1+{K}}(t)\right]_{a,b=0}^{n-1}.$
With reference to the notation of [23, Definition 3.1], the relationship is then
$\tau_{n}(t;{K},{\rho})=(-1)^{{K}n}T^{({\rho}-{K}-2n)}_{n+{K}-1,n}(t)=(-1)^{{K}n}t^{\frac{n(1-n)}{2}}a_{n+{K}-1,n}\tau^{({\rho}-{K}-2n)}_{n+{K}-1,n}(t),$
where in the last identity we have used Lemma 4.5 of loc. cit. to connect with the definition (1.2). The proportionality constant $a_{m,n}$ is defined in the same lemma.
@@
Remark 2.3.
We are considering the moment functional of type (2.3) that corresponds to the choice of contour $\gamma$ surrounding the segment $[0,1]$ in the complex plane. The general theory of moment functionals discussed earlier allows for other choices. While these do not lead to polynomial moments (in the variable $t$), their Hankel determinants would, by the same reasoning in [7], provide isomonodromic tau functions of Painlevé V. Up to the map $z\mapsto 1/z$ and setting ${\rho}={K}=0$ and choosing the contour of integration as the segment $[0,1]$ (the so-called “hard-edge case”), one would recover the moment functional appearing in [5]. In that case, the orthogonal polynomials are the so-called “kissing polynomials”. It would be possible to perform a similar asymptotic analysis of the corresponding tau functions for large $n$ and determine the zero distribution, despite the tau functions not being polynomials.
@@ 2.2 From orthogonal polynomials to Lax pairs
Problem 2.4.
Given $n\in\mathbb{N}$, find a $2\times 2$ matrix-valued function $Y(z)=Y_{n}(z)$ such that
  1. $(1)$
@@ 2. $(2)$
    $Y(z_{+})=Y(z_{-})\begin{bmatrix}1&{\rm e}^{\theta(z)}\\ 0&1\end{bmatrix},\qquad\forall z\in\gamma$
@@
  3. $(3)$
    As $z\to\infty$, the matrix $Y_{n}(z)$ admits an asymptotic expansion of the form
    $Y_{n}(z)=\left(\mathbf{1}+\sum_{\ell=1}^{\infty}\frac{Y_{n}^{(\ell)}}{z^{\ell}}\right)z^{n\sigma_{3}},\qquad\sigma_{3}:=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$, (2.6)
    where the coefficient matrices $Y_{n}^{(\ell)}$ are independent of $z$ and $\mathbf{1}$ denotes the identity matrix of the appropriate size ((here, $2\times 2)$.
In the following theorem, we condense the essential results that are at the core of much of the theory of orthogonal polynomials. The theorem is mostly the reformulation of results of [28] together with notational adaptations to the current context.
Theorem 2.5 ([28]).
Consider the Riemann–Hilber Problem 2.4.
  1. $(1)$
    If the solution exists, it is unique and $\det Y_{n}(z)\equiv 1$.
  2. $(2)$